# TSI-Bench: Benchmarking Time Series Imputation

### v1

@@ 3.3 Models
To comprehensively evaluate the effectiveness of various models in the time series imputation task, we employ 28 different models already incorporated into the PyPOTS Ecosystem, each based on distinct architectures and focusing on different original tasks. These models include Transformer architectures like Transformer [56], Pyraformer [57], Autoformer [31], Informer [42], Crossformer [43], PatchTST [33], ETSformer [58], Nonstationary (short for Nonstationary Transformer) [59], SAITS [20], and iTransformer [60]; RNN architectures such as MRNN [61], BRITS [19], and GRUD [62]; CNN architectures including MICN [63], SCINet [64] and TimesNet [13]; GNN architectures like StemGNN [34]; MLP architectures like FiLM [65], DLinear [44], Koopa [66] and FreTS [67]; Generative architectures based on VAE, GAN, and diffusion models such as GP-VAE [28], US-GAN [68], and CSDI [30]. Besides, we also applied some Traditional methods such as Mean, Median, LOCF (Last Observation Carried Forward), and Linear interpolation. For the original forecasting backbones (Transformer, Pyraformer, Autoformer, etc.), the embedding strategy and the training methodology from SAITS [20] are applied to reimplement them as imputation methods. Please refer to Appendix B and Appendix C.5 for more details of the benchmarked models as well as the systematic paradigm to leverage time series forecasting models for imputation purposes.
@@ 3.4 Implementation Details
Dataset Preprocessing: Except for Physionet2012, Physionet2019, and Pedestrian which contain separated time series samples already, all other datasets are split into the train, validation, and test sets according to the time period, then the sliding window function is applied to generate separated data samples. The standardization is applied to samples of all datasets. Refer to Appendix A.2 for more dataset preprocessing details. All preprocessed datasets are available on Google Drive at this link. Missingness Patterns: To make missing patterns diverse, we apply three different types of missingness (point, subsequence, and block) in our experiments. For the vivid illustration, please refer to C.1. Evaluation Metrics: MAE (mean absolute error), MSE (mean squared error), and MRE (mean relative error) are applied to evaluate the imputation performance. Their mathematical equations are put in Appendix C.2. Besides, inference time and the number of trainable parameters are also collected for discussion. Hyper-parameter Optimization: To obtain fair comparisons and conclusions, hyper-parameter optimization is performed for all deep-learning imputation algorithms. This functionality is implemented with PyPOTS [50] and NNI [69]. The details are available in Appendix C.3. Downstream Task Design: To explore how imputation affects downstream analysis, we perform various tasks with XGBoost, RNN, and Transformer on the imputed datasets for evaluation. Refer to Appendix C.4 for design details.
4 Experiments
Table 1: Imputation performance with 10% point missingness, shown as MAE (standard deviation). For each dataset, the best performance is indicated by bold black text, the second-best performance by normal black text, and the third-best performance by bold dark gray.
@@ 4 Experiments
In this section, we conduct a comprehensive analysis based on up to 34,804 experiments of 28 algorithms across 8 different domain datasets. Our experiments are divided into three main parts: (1) Data Perspective: We test different missingness ratios (10%, 50%, and 90%) and various missingness patterns (point, subsequence, block) across all datasets. (2) Model Perspective: We analyze all methods according to target task type (forecasting, imputation) and architecture type (Transformer/attention, RNN, CNN, MLP, etc.), while also considering the impact of model size and inference time on practical applications. (3) Downstream Task Perspective: We evaluate the imputed data quality using three different architectures (XGBoost, RNN, Transformer) across three classical time-series downstream tasks (classification, regression, and forecasting) to demonstrate the significance and importance of imputation in downstream tasks. Full results are listed in Appendix D.
Figure 3: Performance of 50% block missing.
Figure 4: Performance of 50% subseq missing.
@@ 4.1 Data Perspective
We conduct the following three experiments on data and missingness patterns: datasets from different application domains, missingness rates, and different missing features.
@@
Application Domain Table 1 shows that different models exhibit significant performance variations across datasets from different application domains. For instance, Autoformer performs well on the Pedestrian dataset (with an MAE of 0.197) but performs poorly on the ItalyAir dataset (with an MAE of 0.295). This trend is also observed within the same application domain. BRITS excels on the BeijingAir dataset (with an MAE of 0.127) but shows worse performance on the ItalyAir dataset (with an MAE of 0.235). Overall, different models demonstrate varying performance across different application domains and datasets, with each model having its own strengths and weaknesses in specific datasets. Even in simpler application fields, such as electricity, we can get better results using traditional LOCF and Linear methods. It seems that we cannot find a model that performs perfectly in all application scenarios.
@@
Missingness Rate As the missingness data rate increases from 10% to 90%, there is a general trend of performance degradation across all models and datasets, as shown in Appendix D. For instance, Autoformer’s MAE on the BeijingAir dataset rises from 0.257 at 10% missingness to 0.898 at 50% and 0.806 at 90% missing. Different models exhibit varying sensitivity to missing data, with BRITS maintaining relatively lower errors compared to others, whereas CSDI shows extreme sensitivity. For example, CSDI’s MAE on the ItalyAir dataset jumps from 0.539 at 10% missingness to 4.492 at 90%. CSDI is a conditional diffusion model that performs poorly and becomes less stable with high missingness rates primarily due to noise accumulation, lack of contextual information, sparsity of training data, and uncertainty in the imputation targets. Computational time also varies, with CSDI showing substantially longer times at higher missingness rates, while models like Autoformer and DLinear remain consistent. We want to highlight the importance of selecting appropriate models based on the expected missing data rate. For models with similar performance, it is better to consider computational efficiency for practical implementation.
@@
Missingness Pattern It is evident that different patterns significantly impact model performance, as indicated in Figures 4 and 4. Overall, block and subsequence missingness patterns result in higher error metrics compared to point missing, indicating that continuous missing data has a greater adverse effect on model performance. Specifically, point missingness has the least impact on most models, as they can easily interpolate or estimate the missing data from adjacent points; for example, BRITS performs excellently under this pattern. In contrast, block missingness leads to more significant information loss over extended periods, making it challenging for models to infer missing data from distant points, as seen with the increased MAE and MSE for Autoformer and Crossformer. Subsequence missingness further deteriorates model performance since the missing subsequences may contain critical time patterns or trends that the models cannot easily reconstruct. Thus, different missingness patterns affect model performance to varying degrees, with continuous missingness patterns (block and subsequence missing) posing greater challenges.
@@
Remark 1: We suggest choosing imputation models tailored to the specific data characteristics, such as application domains, and missing patterns. Practitioners can leverage our comprehensive insights to make informed model choices to narrow their trials and experiments in selecting models, enhancing the development and open-source accessibility of time series imputation.
@@ 4.2 Model Perspective
Due to the space limit, we limit our discussions in this section to the model architectures and the original task based on 10% missingness rate.
@@
Architecture By analyzing the inference time, model size, and performance metrics of different architectures from Figure 5 and Table 1, we can draw the following conclusions: Transformer/Attention-based models, such as Autoformer and Informer, demonstrate excellent performance, but some models like Crossformer and SAITS have larger model sizes, potentially leading to lower inference efficiency. RNN-based models like BRITS show good performance (low MSE, MAE) but have longer inference times and moderate model sizes. CNN-based models like SCINet have faster inference times but relatively larger model sizes, with slightly lower performance compared to Transformer and RNN architectures. MLP-based models like DLinear have clear advantages in inference time and model size, with moderate performance, making them suitable for applications requiring quick inference and smaller models. CSDI, as a generative model, performs well on some metrics (e.g., MAE) but has extremely long inference time and moderate model size, suitable for scenarios demanding very high performance. Traditional methods, due to their limited scope, usually do not yield optimal results, especially in datasets with high missingness rates and complexity. However, they can be more efficient and perform better on simpler datasets In summary, different architecture models have their own strengths and weaknesses in terms of inference time, model size, and performance metrics, necessitating the selection of an appropriate model based on specific application requirements.
@@
Forecasting Backbones for Imputation With the adaptation paradigm C.5, we transfer the backbone of the time series forecasting models to the imputation task and compare their performance against typical imputation methods. Results are categorized into Table 1, and the detailed descriptions of the backbones are discussed in Appendix B. For forecasting backbones, Crossformer has a longer inference time of 6.77 seconds, a larger model size of 52,933,788, but stable performance, with an MAE of 0.184 on the Beijing dataset. DLinear, with the shortest inference time of 0.17 seconds and a model size of 204,728, shows efficient performance with an MAE of 0.215 on the Beijing dataset, making it ideal for real-time applications. In imputation, BRITS has an inference time of 17.04 seconds, a model size of 3,598,496, and performs excellently on the Beijing dataset with an MAE of 0.127. CSDI, however, has a long inference time of 417.52 seconds, a model size of 244,833, and higher MSE values, limiting its practical use despite advanced techniques. Generally, methods originally designed for imputation do not exhibit extraordinary performance compared to those designed for forecasting. In many scenarios, the adapted forecasting methods tend to perform better. A potential reason is that there are more papers on time series forecasting, compared to time series imputation. Therefore, models in time series forecasting are more updated and advanced in capturing temporal information and multivariate correlations. The experimental results demonstrate the success of our adaptation paradigm.
@@ 4.3 Downstream Tasks Perspective
Figure 7: Comparison of downstream classification performance on the PhysioNet2012 dataset with 10% point missing data.
Classification Task In this ablation test, we first apply different models performing the imputation task then conduct the classification task based on three types of classifiers. We would like to explore the performance improvement of various imputation methods for the downstream classification task, as shown in Figure 7. For example, iTransformer shows a ROC_AUC of 0.771 without imputation, which improves significantly to 0.852 after imputation. Besides, the nonstationary method performs best with imputation, achieving a ROC_AUC of 0.860, demonstrating strong robustness in handling missing data and positively contributing to downstream tasks. Overall, imputation plays a crucial role in the classification task for preprocessing, as the performance of most methods significantly improving after imputation. Also, selecting an appropriate imputation method can enhance downstream performance. Therefore, future research and applications might place greater emphasis on the development and optimization of imputation techniques to improve the robustness and effect of various downstream classification algorithms.
Table 2: Regression performance comparison in ETT_h1 datasets with 50% point missing, 50% block missing and 50% subsequence missing.
@@
iTransformer 1.224 (0.000) 1.377 (0.062) 1.406 (0.069) 1.170 (0.000) 1.434 (0.047) 1.470 (0.053) 1.208 (0.000) 1.422 (0.075) 1.399 (0.080)
SAITS 1.175 (0.000) 1.402 (0.056) 1.401 (0.083) 1.094 (0.000) 1.424 (0.077) 1.469 (0.054) 1.168 (0.000) 1.363 (0.072) 1.385 (0.057)
Nonstationary 1.227 (0.000) 1.446 (0.061) 1.384 (0.087) 1.284 (0.000) 1.448 (0.053) 1.469 (0.036) 1.189 (0.000) 1.438 (0.059) 1.368 (0.055)
ETSformer 1.222 (0.000) 1.329 (0.060) 1.351 (0.061) 1.187 (0.000) 1.407 (0.059) 1.433 (0.058) 1.004 (0.000) 1.285 (0.064) 1.327 (0.096)
PatchTST 1.234 (0.000) 1.398 (0.059) 1.370 (0.074) 1.253 (0.000) 1.476 (0.079) 1.502 (0.042) 1.183 (0.000) 1.362 (0.078) 1.340 (0.074)
Crossformer 1.228 (0.000) 1.391 (0.058) 1.380 (0.070) 1.145 (0.000) 1.449 (0.058) 1.474 (0.042) 1.149 (0.000) 1.335 (0.070) 1.285 (0.081)
Informer 1.214 (0.000) 1.391 (0.059) 1.398 (0.076) 1.196 (0.000) 1.423 (0.076) 1.442 (0.042) 1.158 (0.000) 1.333 (0.080) 1.351 (0.055)
Autoformer 1.348 (0.000) 1.450 (0.108) 1.442 (0.155) 1.364 (0.000) 1.471 (0.061) 1.538 (0.072) 1.327 (0.000) 1.267 (0.047) 1.363 (0.204)
Pyraformer 1.169 (0.000) 1.366 (0.057) 1.373 (0.071) 1.081 (0.000) 1.391 (0.077) 1.443 (0.042) 1.071 (0.000) 1.328 (0.069) 1.330 (0.053)
Transformer 1.170 (0.000) 1.378 (0.056) 1.357 (0.081) 1.074 (0.000) 1.391 (0.060) 1.350 (0.090) 1.128 (0.000) 1.296 (0.064) 1.277 (0.077)
BRITS 1.177 (0.000) 1.371 (0.061) 1.364 (0.076) 0.999 (0.000) 1.153 (0.042) 1.146 (0.036) 0.826 (0.000) 1.024 (0.010) 1.000 (0.057)
MRNN 1.070 (0.000) 1.250 (0.076) 1.379 (0.058) 1.111 (0.000) 1.362 (0.061) 1.483 (0.055) 0.870 (0.000) 1.274 (0.078) 1.387 (0.053)
GRUD 1.205 (0.000) 1.339 (0.061) 1.326 (0.074) 1.123 (0.000) 1.351 (0.058) 1.362 (0.044) 1.043 (0.000) 1.286 (0.046) 1.211 (0.080)
TimesNet 1.127 (0.000) 1.333 (0.054) 1.316 (0.066) 1.213 (0.000) 1.446 (0.058) 1.460 (0.041) 1.064 (0.000) 1.316 (0.068) 1.294 (0.101)
MICN 1.163 (0.000) 1.305 (0.078) 1.379 (0.070) 1.121 (0.000) 1.442 (0.052) 1.503 (0.060) 1.145 (0.000) 1.338 (0.066) 1.344 (0.073)
SCINet 1.174 (0.000) 1.411 (0.055) 1.375 (0.080) 1.181 (0.000) 1.450 (0.061) 1.446 (0.038) 1.182 (0.000) 1.361 (0.075) 1.323 (0.065)
StemGNN 1.139 (0.000) 1.392 (0.057) 1.410 (0.078) 1.211 (0.000) 1.391 (0.067) 1.439 (0.045) 1.136 (0.000) 1.373 (0.064) 1.323 (0.052)
FreTS 1.207 (0.000) 1.406 (0.056) 1.397 (0.082) 1.159 (0.000) 1.456 (0.079) 1.514 (0.045) 1.163 (0.000) 1.356 (0.080) 1.355 (0.068)
Koopa 1.227 (0.000) 1.384 (0.064) 1.419 (0.062) 1.233 (0.000) 1.446 (0.083) 1.509 (0.046) 1.274 (0.000) 1.347 (0.063) 1.302 (0.052)
DLinear 1.216 (0.000) 1.380 (0.056) 1.397 (0.076) 1.212 (0.000) 1.466 (0.058) 1.512 (0.044) 1.166 (0.000) 1.368 (0.064) 1.359 (0.114)
FiLM 1.247 (0.000) 1.379 (0.057) 1.442 (0.048) 1.422 (0.000) 1.389 (0.057) 1.455 (0.036) 1.301 (0.000) 1.391 (0.057) 1.368 (0.060)
CSDI 1.136 (0.000) 1.373 (0.050) 1.393 (0.070) 1.191 (0.000) 1.354 (0.049) 1.413 (0.041) 1.202 (0.000) 1.355 (0.048) 1.362 (0.054)
US-GAN 1.155 (0.000) 1.299 (0.040) 1.283 (0.055) 1.004 (0.000) 1.244 (0.052) 1.219 (0.079) 0.951 (0.000) 1.174 (0.069) 1.098 (0.145)
GP-VAE 1.178 (0.000) 1.336 (0.066) 1.340 (0.076) 1.138 (0.000) 1.310 (0.046) 1.354 (0.077) 0.996 (0.000) 1.177 (0.061) 1.184 (0.077)
Mean 1.431 (0.000) 1.512 (0.120) 1.627 (0.072) 1.430 (0.000) 1.859 (0.095) 1.779 (0.035) 1.413 (0.000) 1.669 (0.083) 1.750 (0.043)
Median 1.517 (0.086) 1.531 (0.128) 1.671 (0.069) 1.499 (0.069) 1.858 (0.089) 1.777 (0.034) 1.496 (0.083) 1.688 (0.100) 1.736 (0.061)
LOCF 1.446 (0.123) 1.477 (0.134) 1.571 (0.158) 1.446 (0.094) 1.699 (0.237) 1.651 (0.182) 1.433 (0.112) 1.601 (0.150) 1.611 (0.188)
Linear 1.393 (0.140) 1.453 (0.127) 1.525 (0.163) 1.423 (0.091) 1.610 (0.258) 1.588 (0.193) 1.395 (0.118) 1.558 (0.151) 1.556 (0.191)
Regression Task As described in Table 2, imputation methods with different missingness patterns (50% point, subsequence, and block missing) significantly affect the performance of downstream regression tasks. For instance, the BRITS model, which is an RNN-based method, achieves the lowest MAE of 0.826 in the block missingness scenario, indicating its robustness in maintaining data integrity and its effectiveness in supporting accurate regression outcomes. This suggests that the ability of BRITS to capture temporal dependencies despite substantial data loss translates into more reliable regression predictions. Similarly, Transformer-based models like Autoformer and Informer demonstrate strong performance, with MAE values of 1.327 and 1.158 respectively in the block missingness scenario, highlighting their effectiveness in handling complex missing data patterns and enhancing the predictive accuracy of downstream regression tasks. Conversely, traditional imputation methods such as Mean and Median show higher MAE values, 1.413 and 1.496 respectively, in the same scenario, underscoring their limitations and resulting in less accurate regression results. The performance of a simple XGBoost model applied to imputed data, which outperforms untuned deep learning classifiers, further reinforces the importance of robust imputation. This demonstrates that high-quality imputation not only preserves the integrity of the data but also significantly enhances the predictive power of regression models.
We also perform experiments that take forecasting as a downstream task in Appendix D.7.3.
Remark 3: Imputation is not only crucial in itself but also plays a vital role in enhancing the performance of downstream tasks, such as classification, regression, and forecasting. By filling in missing data, imputation helps in recovering the original temporal information of the original dataset, which is beneficial for reliable analysis and modeling.

### v2

@@ 3.3 Models
To comprehensively evaluate the performance of various models in the time series imputation task, we utilise 28 different models integrated into the PyPOTS Ecosystem. These models are based on distinct architectures, originally designed for different tasks. Among the evaluated models are Transformer architectures, including Transformer (Yıldız et al., 2022), Pyraformer (Liu et al., 2021), Autoformer (Wu et al., 2021), Informer (Zhou et al., 2021), Crossformer (Zhang & Yan, 2022), PatchTST (Nie et al., 2023), ETSformer (Woo et al., 2022), Nonstationary Transformer (Liu et al., 2022c), SAITS (Du et al., 2023), and iTransformer (Liu et al., 2023b); RNN architectures, such as MRNN (Yoon et al., 2018), BRITS (Cao et al., 2018), and GRUD (Che et al., 2018b); CNN architectures, including MICN (Wang et al., 2022), SCINet (Liu et al., 2022a), and TimesNet (Wu et al., 2023); GNN architectures, such as StemGNN (Cao et al., 2020); MLP architectures, such as FiLM (Zhou et al., 2022), DLinear (Zeng et al., 2023), Koopa (Liu et al., 2024), and FreTS (Yi et al., 2024); and Generative architectures, including GP-VAE (Fortuin et al., 2020), US-GAN (Miao et al., 2021a), and CSDI (Tashiro et al., 2021), which are based on VAE, GAN, and diffusion models. Additionally, we apply several traditional methods, such as Mean, Median, LOCF (Last Observation Carried Forward), and Linear interpolation. For the original forecasting models (e.g., Transformer, Pyraformer, Autoformer), we adapt the embedding strategies and training methodologies from SAITS (Du et al., 2023) to reimplement them as imputation methods. Further details on the benchmarked models and the systematic approach used to adapt time series forecasting models for imputation purposes can be found in Appendix B and Appendix 10.
@@ 3.4 Implementation Details
Dataset Preprocessing: Except PhysioNet2012, PhysioNet2019, and Pedestrian, which already contain separated time series samples, all other datasets are split into training, validation, and test sets according to the time period. A sliding window function is then applied to generate individual data samples. Standardization is performed on all datasets. For further details on dataset preprocessing, please refer to Appendix A.2. Missing Patterns: To more effectively simulate missing scenarios in time series, we deviate from the traditional categories of missing mechanisms introduced by (Rubin, 1976). Instead, we implement point missing, subsequence missing, and block missing, which are particularly representative of common missing data situations in time series (Mitra et al., 2023; Cini et al., 2022; Khayati et al., 2020). These patterns capture both random and structured missingness, reflecting the dynamic nature of time series data and providing a robust framework for evaluating imputation strategies. For a detailed discussion and visual representation of these missing patterns, please refer to Appendix C.1. Evaluation Metrics: To assess imputation performance, we use mean absolute error (MAE), mean squared error (MSE), and mean relative error (MRE) (Cao et al., 2018). The mathematical formulations for these metrics can be found in Appendix C.2. Additionally, we collect inference time and the number of trainable parameters for further discussion. Hyper-parameter Optimisation: To ensure fair comparisons, hyper-parameter optimisation is applied to all deep-learning-based imputation algorithms. This process is implemented using PyPOTS (Du, 2023) and NNI (Microsoft, 2021), with details provided in Appendix C.3. Downstream Task Design: To evaluate the impact of imputation on downstream analysis, we perform various tasks using XGBoost, RNN, and Transformer models on the imputed datasets. Given that XGBoost has an inherent capability to manage incomplete data, we set XGBoost without imputation as the baseline for comparison. Further details on the downstream task design can be found in Appendix C.4.
4 Experimental analysis
Table 2: Imputation performance with 10% point missingness, shown as MAE (standard deviation). For each dataset, the best performance is indicated by the bold black text, the second-best performance by normal black text, and the third-best performance by bold dark grey.
@@ 4 Experimental analysis
We conduct a comprehensive analysis involving up to 34,804 experiments across 28 algorithms and 8 diverse domain datasets. The experiments are structured to examine three key perspectives: (1) Data Perspective: We evaluate different missing rates (10%, 50%, and 90%) and missing patterns (point, subsequence, block) across all datasets. (2) Model Perspective: We analyze each method based on its target task type (forecasting, imputation) and architecture (Transformer/attention, RNN, CNN, MLP, etc.), while also assessing the influence of model size and inference time on practical applications. (3) Downstream Task Perspective: We assess the quality of the imputed data using three architectures (XGBoost, RNN, Transformer) across three classical time series downstream tasks (classification, regression, forecasting) to highlight the importance and impact of imputation on downstream performance. Full results are available in Appendix D.
Figure 3: Performance of 50% block missing (left) and 50% subseq missing (right).
@@ 4.1 Data Perspective
We conduct experiments on the following aspects of data and missing patterns: datasets from different application domains, missing rates, and different missing patterns.
@@
Application Domains Table 2 shows that different models show significant performance disparities across datasets from different application domains. For instance, Autoformer performs well on the Pedestrian dataset (with an MAE of 0.197) but performs poorly on the ItalyAir dataset (with an MAE of 0.295). This trend is also observed within the same application domain. BRITS excels on the BeijingAir dataset (with an MAE of 0.127) but shows worse performance on the ItalyAir dataset (with an MAE of 0.235). Overall, different models demonstrate varying performance across different application domains and datasets, with each model having its own strengths and weaknesses in specific datasets. Even in simpler application fields, such as electricity, we can get better results using traditional LOCF and Linear methods.
@@
Missing Rates As the missingness data rate increases from 10% to 90%, there is a general trend of performance degradation across all models and datasets, as shown in Appendix D. For instance, Autoformer’s MAE on the BeijingAir dataset rises from 0.257 at 10% to 0.898 at 50% missingness. Different models exhibit varying sensitivity to missing data, with BRITS maintaining relatively lower errors compared to others, whereas CSDI shows extreme sensitivity. CSDI’s MAE on the ItalyAir dataset jumps from 0.539 at 10% missingness to 4.492 at 90%. CSDI is a conditional diffusion model that performs poorly and becomes less stable with high missing rates primarily due to noise accumulation and lack of contextual information. Computational time also varies, with CSDI showing substantially longer times at higher missing rates, while models like Autoformer and DLinear remain consistent. We want to highlight the importance of selecting appropriate models based on the expected missing data rate. For models with similar performance, it is better to consider computational efficiency for practical implementation.
@@
Missing Patterns It is evident that different patterns significantly impact model performance, as shown in Figure 3. In general, block and subsequence missing patterns result in higher error metrics compared to point missingness, indicating that continuous missing data have a greater adverse effect on model performance. Specifically, point missingness has the least impact on most models, because missing data can be easily interpolated from adjacent points; for example, BRITS performs excellently under this pattern. In contrast, block missingness leads to more significant information loss over extended periods, making it challenging for models to infer missing data from distant points, as seen with the increased MAE and MSE for Autoformer and Crossformer. Subsequence missingness further degrades model performance since the missing subsequences may contain critical temporal patterns or trends that the models cannot easily reconstruct. Thus, different missing patterns affect model performance to varying degrees, with continuous missing patterns (block and subsequence missing) posing greater challenges.
We further provide visualization of imputed data from all imputation methods on ETT_h1 and Electricity in Appendix D.6. As shown in Figure 12 and Figure 13, different features within the same dataset exhibit substantial variation. Some features like Feature #76 in Electricity exhibit limited temporal dependencies, whereas others like Feature #1 in ETT_h1 present clear temporal patterns, such as step changes. Moreover, when comparing the same feature across different missing patterns (e.g., point versus block missingness), the imputation results can differ significantly. While most algorithms perform well with point missingness, they often struggle with block missingness, which demands the preservation of longer-term dependencies. The results suggest that it is challenging for any single imputation algorithm to consistently perform well across different datasets and missingness scenarios.
@@
Remark 1: We suggest choosing imputation models tailored to data characteristics, such as application domains and missing patterns. Practitioners can leverage our comprehensive insights to make informed model choices to narrow their trials and experiments in selecting models, enhancing the development and open-source accessibility of time series imputation.
@@ 4.2 Model Perspective
Given the space constraints, we will focus our discussion in this section on the model architectures and the primary task assuming a 10% missing rate.
@@
Architecture The inference time, model size, and performance metrics of different architectures shown in Figure 4 and Table 2 indicate the following: Transformer/attention-based models, such as Autoformer and Informer, demonstrate excellent performance, but some models like Crossformer and SAITS have larger model sizes, potentially leading to lower inference efficiency. RNN-based models like BRITS show good performance (low MSE, MAE) but have longer inference times and moderate model sizes. CNN-based models like SCINet have faster inference times but relatively larger model sizes, with slightly lower performance compared to Transformer and RNN architectures. MLP-based models like DLinear have clear advantages in inference time and model size, with moderate performance, making them suitable for applications requiring quick inference and smaller models. CSDI, as a generative model, performs well on MAE, but has an extremely long inference time and moderate model size, suitable for scenarios demanding very high performance. Traditional methods, due to their limited scope, usually do not yield optimal results, especially in datasets with high missing rates and complexity. However, they can be more efficient and perform better on simpler datasets. In summary, different architectures have their own strengths and weaknesses in terms of inference time, model size, and performance metrics, necessitating the selection of an appropriate model based on specific application requirements.
@@
Forecasting Backbones for Imputation With the adaptation paradigm in Appendix 10, we transfer the backbone of the time series forecasting models to the imputation task and compare their performance against typical imputation methods. The results are categorized in Table 2, and detailed descriptions of the backbones are discussed in Appendix B. For forecasting backbones, DLinear, with the shortest inference time of 0.17 seconds and a model size of 204K, shows efficient performance with an MAE of 0.215 on the Beijing dataset, making it ideal for real-time applications. In imputation, BRITS has an inference time of 17.04 seconds, a model size of 3.5M, and performs excellently on the Beijing dataset with an MAE of 0.127. CSDI, however, has a long inference time of 417.52 seconds, a model size of 244K, and higher MSE values, limiting its practical use despite advanced techniques. Generally, methods originally designed for imputation do not exhibit extraordinary performance compared to those designed for forecasting. In many scenarios, the adapted forecasting methods tend to perform better. A potential reason is that more research has been conducted on time series forecasting compared to time series imputation. Therefore, time series forecasting models are more updated and advanced in capturing temporal information and multivariate correlations. The experimental results demonstrate the success of our adaptation paradigm.
@@ 4.3 Downstream Tasks Perspective
Figure 6: Comparison of downstream performance on the PhysioNet2012 dataset for classification with 10% point missing (left) and the ETT_h1 dataset for forecasting with 50% point missing (right).
Forecasting Task We evaluated the downstream forecasting performance of various models on the ETT_h1 dataset. As shown in Figure 6 (right), certain imputation methods significantly improve downstream forecasting performance. For instance, the Nonstationary imputation method consistently enhances accuracy across multiple models. Transformer-based models show notable gains with robust imputation strategies. XGBoost without imputation results in a higher MAE, but applying Transformer models with imputation consistently reduces the MAE. Overall, selecting the right imputation method is crucial for improving downstream forecasting task performance.
Classification Task In this test, we apply different models for imputation, followed by three types of classifiers. The goal is to assess the performance improvement brought by various imputation methods, as shown in Figure 6 (left). For instance, iTransformer achieves a ROC_AUC of 0.771 without imputation, which increases to 0.852 after imputation. The non-stationary method performs best with an ROC_AUC of 0.860, showcasing strong robustness in handling missing data. Overall, imputation significantly enhances classification performance, and the choice of imputation method is crucial for optimizing downstream tasks.
Table 3: Regression performance in ETT_h1 datasets with 50% point missing, 50% block missing and 50% subsequence missing. For each model, the best performance is indicated by the bold black text, the second-best performance by normal black text, and the third-best performance by bold dark grey.
@@
iTransformer 1.224 1.377 1.406 1.170 1.434 1.470 1.208 1.422 1.399
SAITS 1.175 1.402 1.401 1.094 1.424 1.469 1.168 1.363 1.385
Nonstationary 1.227 1.446 1.384 1.284 1.448 1.469 1.189 1.438 1.368
ETSformer 1.222 1.329 1.351 1.187 1.407 1.433 1.004 1.285 1.327
PatchTST 1.234 1.398 1.370 1.253 1.476 1.502 1.183 1.362 1.340
Crossformer 1.228 1.391 1.380 1.145 1.449 1.474 1.149 1.335 1.285
Informer 1.214 1.391 1.398 1.196 1.423 1.442 1.158 1.333 1.351
Autoformer 1.348 1.450 1.442 1.364 1.471 1.538 1.327 1.267 1.363
Pyraformer 1.169 1.366 1.373 1.081 1.391 1.443 1.071 1.328 1.330
Transformer 1.170 1.378 1.357 1.074 1.391 1.350 1.128 1.296 1.277
BRITS 1.177 1.371 1.364 0.999 1.153 1.146 0.826 1.024 1.000
MRNN 1.070 1.250 1.379 1.111 1.362 1.483 0.870 1.274 1.387
GRUD 1.205 1.339 1.326 1.123 1.351 1.362 1.043 1.286 1.211
TimesNet 1.127 1.333 1.316 1.213 1.446 1.460 1.064 1.316 1.294
MICN 1.163 1.305 1.379 1.121 1.442 1.503 1.145 1.338 1.344
SCINet 1.174 1.411 1.375 1.181 1.450 1.446 1.182 1.361 1.323
StemGNN 1.139 1.392 1.410 1.211 1.391 1.439 1.136 1.373 1.323
FreTS 1.207 1.406 1.397 1.159 1.456 1.514 1.163 1.356 1.355
Koopa 1.227 1.384 1.419 1.233 1.446 1.509 1.274 1.347 1.302
DLinear 1.216 1.380 1.397 1.212 1.466 1.512 1.166 1.368 1.359
FiLM 1.247 1.379 1.442 1.422 1.389 1.455 1.301 1.391 1.368
CSDI 1.136 1.373 1.393 1.191 1.354 1.413 1.202 1.355 1.362
USGAN 1.155 1.299 1.283 1.004 1.244 1.219 0.951 1.174 1.098
GPVAE 1.178 1.336 1.340 1.138 1.310 1.354 0.996 1.177 1.184
Mean 1.431 1.512 1.627 1.430 1.859 1.779 1.413 1.669 1.750
Median 1.517 1.531 1.671 1.499 1.858 1.777 1.496 1.688 1.736
LOCF 1.446 1.477 1.571 1.446 1.699 1.651 1.433 1.601 1.611
Linear 1.393 1.453 1.525 1.423 1.610 1.588 1.395 1.558 1.556
Regression Task As shown in Table 3, imputation methods significantly affect downstream regression performance under three missing patterns. For example, the RNN-based BRITS model achieves the lowest MAE of 0.826 in the block missing and subsequence scenarios, demonstrating its robustness in handling data loss. Transformer-based models also perform well, highlighting their ability to handle complex missing data patterns. In contrast, traditional methods such as Mean and Median imputation result in higher MAE values, underscoring their limitations. The improved performance of models like XGBoost applied to imputed data further emphasizes the importance of robust imputation for enhancing regression performance.
Remark 3: Imputation is not only critical on its own but also plays a key role in improving the performance of downstream tasks. Additionally, pre-imputation demonstrates high adaptability across different downstream tasks within the same datasets. Once high-quality imputed time series are generated, they can be effectively leveraged across a wide range of downstream applications.